A and B are independent events. Find the missing probability.
P(A)= 3/7, P(A and B)= 1/3. find P(b)

Answer by stanbon(75887)

A and B are independent events. Find the missing probability.
P(A)= 3/7, P(A and B)= 1/3. find P(b)
-----------
If A and B are independent P(A and B) = P(A)*P(B)
1/3 = (3/7)*P(B)
P(B) = (1/3)/(3/7) = 7/9
=============
Cheers,
Stan H.